OpenStudy (anonymous):

Solve the following system of equations: -2x + y = 1 -4x + y = -1 (3, 1) (-1, 3) (-1, -3) (1, 3)

OpenStudy (anonymous):

Please Help! I will fan and medal!

OpenStudy (noelgreco):

Use elimination by subtracting the second equation from the first. Watch your signs!

OpenStudy (anonymous):

Would it be D? @NoelGreco

OpenStudy (anonymous):

Hello? @NoelGreco

OpenStudy (anonymous):

-2x + y = 1 (equation 1) -4x + y = -1 (equation 2) for equation 1 -2x + y = 1 y = 1 + 2x (equation 3) substitute this equation 3 to equation 2, then -4x + y = -1 -4x + (1 + 2x) = -1 -4x + 1 + 2x = -1 -2x + 1 = -1 -2x = -1 - 1 -2x = -2 x = 2/2 x=1

OpenStudy (anonymous):

then for y y = 1 + 2x y = 1 + 2(1) y = 1 + 2 y = 3

OpenStudy (anonymous):

So the answer is (x,y) = (1,3)

OpenStudy (anonymous):

It'd D!!!!

OpenStudy (anonymous):

Thank you so much! @gerryliyana